OpenStudy (anonymous):

Solve for x in the following equation: 4 · (125^x/3) − 310536 = 1964 how do i get ride of the power x/3?

hartnn (hartnn):

first add 310536 to isolate the term with 1/3 expnent

hartnn (hartnn):

then you can cube both sides

OpenStudy (anonymous):

4*125^x/3=33020 do I multiply the 125 by 4? first

hartnn (hartnn):

no, you can't and how is it 33020 ? i get 312500

OpenStudy (anonymous):

I missed a number, sorry.

hartnn (hartnn):

(125^(x/3)) is one term, to isolate it, you'll have to divide by 4 on both sides.

OpenStudy (anonymous):

125^x/3=78125

hartnn (hartnn):

yes now lets take log on both sides

hartnn (hartnn):

use the property that log A^B = B log A log 125^(1/3) =...?

OpenStudy (anonymous):

so log 125^(x/3)=3logx

hartnn (hartnn):

try again? where did 125 go ? note that the exponent becomes the co-efficient log A^B = B log A log 125^(1/3) =...?

OpenStudy (ranga):

A = 125 B = 1/3

OpenStudy (anonymous):

why not B=x/3 and um, I'm lost

OpenStudy (ranga):

yep B = x/3 sorry.

OpenStudy (anonymous):

so… it'll be x/3log125 but then what happened to the 78125?

hartnn (hartnn):

yeah typo log 125^(x/3) =...

hartnn (hartnn):

yes we take log both sides so we get (x/3) log 125 = log 78125 got this ?

OpenStudy (anonymous):

why does the log go on both sides?

OpenStudy (anonymous):

sorry about my questions, I am trying to understand

hartnn (hartnn):

its good that you're trying to understand :) what we do on one side we must do on other side too. so that the equation remains unchanged so if we take log, we need to take log on both sides.

OpenStudy (anonymous):

then this is where the C=D comes into play right?

OpenStudy (ranga):

At one point you had the equation: 125^x/3=78125 The x/3 is an exponent. We want to bring it down. And the way to do that is to take the log. But what we do on the left we have to do on the right too in order to keep the equation in tact.

hartnn (hartnn):

no, here we have a co-efficient with the log its not like log C = log D (so you can say C=D) here, its like a logC = log D

hartnn (hartnn):

we have (x/3) log 125 = log 78125 right ? we need to isolate the x start by dividing log 125 on both sides.

OpenStudy (ranga):

We need to solve for x. In the equation (x/3) log 125 = log 78125 everything else is a constant. 3 is a number, log(125) is a number, log(78125) is a number. the calculator will tell you what it is. So what should be done to isolate x? Divide both sides by the n umber log(125). Then multiply both sides by the number 3.

OpenStudy (anonymous):

x/3=625?

hartnn (hartnn):

what did your calculator give you for this ? log 78125/ log 125 =..

hartnn (hartnn):

i think you just did 78125/125

OpenStudy (anonymous):

i can not utilize a calculator for this assignment i have to explain it step by step and show all my work, making side notes as well. so thats exactly what i did divide.

OpenStudy (anonymous):

was that wrong?

hartnn (hartnn):

yeah x/3 log 125 = log 78125 right ? so, x/3 = log 78125 / log 125

hartnn (hartnn):

78125 = 5^7 125 = 5^3

OpenStudy (anonymous):

um, where did the 5^7 come from and the 5^3

hartnn (hartnn):

if i say 8, how would you know its 2^3 ?

hartnn (hartnn):

78125 = 5*5*5*5*5*5*5 = 5^7

OpenStudy (anonymous):

yes but how did you go from x/3=625 to what your doing?

hartnn (hartnn):

x/3 = 625 is incorrect

hartnn (hartnn):

x/3 log 125 = log 78125 right ? so, x/3 = log 78125/log 125 got till here ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

so, x/3 = log 5^7 / log 5^3

OpenStudy (anonymous):

um, ok…

hartnn (hartnn):

now since we know log A^B = B log A log 5^7 = 7 log 5

hartnn (hartnn):

what about log 5^3 ?

OpenStudy (anonymous):

3 log 5

hartnn (hartnn):

yes

hartnn (hartnn):

x/3 = 7 log 5 / 3log 5 what gets cancelled ?

OpenStudy (anonymous):

log5?

hartnn (hartnn):

yep x/3 = 7/3 you can find x from here but did u get all steps till here ?

OpenStudy (anonymous):

yea mostly, so x=7 :)

hartnn (hartnn):

yes! :)